OpenStudy (anonymous):

Use the elimination method to solve the system of equations. Choose the correct ordered pair. 5x - 2y = 9 3x + 4y = -5 A. (1, -2) B. (3, 3) C. (7, 13) D. (5, -5)

OpenStudy (anonymous):

a

OpenStudy (shinalcantara):

Elimination Method: 5x -2y = 9 ----equation 1 3x + 4y = -5 ----equation 2 multiply equation 1 by 2 2(5x -2y = 9) 10x - 4y = 18 ----equation 3 add equations 2 and 3 3x + 4y = -5 + 10x - 4y = 18 13x = 13 x = 1 substitute the value of x to equation 1 5x -2y = 9 5(1) -2y = 9 5 - 2y = 9 -2y = 9 -5 -2y = 4 y = -2 solution (1,-2)

OpenStudy (anonymous):

Thanks Shinal

OpenStudy (anonymous):

And mayur

OpenStudy (anonymous):

Which matrix represents the system of equations shown below? -4x + 6y = 9 8x - y = 7 A. B. C. D.